The fundamental frequency of a string stretched with a weight of 4kg is 256Hz.then weight required to produce its octave

Grade:11

1 Answers

Arun
6 years ago
 
Dear Student
 
If the fundamental frequency of the spring is f
Then frequency of its octave is 2f.
f = K (T)1/2
2f = K(T')1/2
 
Hence T' = 4T
T' = 4*4 = 16 kg- wt
 
Regards,
Arun(AskIITians Forum Expert)